OpenStudy (anonymous):

solve for x: sqrt x+16 - sqrt x-2=0 A. 3 B. 7 C. 9 D. 15

OpenStudy (anonymous):

@jim_thompson5910 i got 9

OpenStudy (anonymous):

It cannot be this way! \[\sqrt{x+16}-\sqrt{x-2}=0\]

OpenStudy (anonymous):

so what would be my answer then?

OpenStudy (anonymous):

is the question the way I wrote?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

if the question is exactly like I wrote, then there can be no solution for it!

OpenStudy (anonymous):

for the second part the sqrt is only over 2x

OpenStudy (anonymous):

check the question and use the equation editor below to enter the equation

OpenStudy (anonymous):

I mean x lol

OpenStudy (anonymous):

see..!!

OpenStudy (anonymous):

so, like this?\[\sqrt{x+16}-\sqrt{x}-2=0\]

OpenStudy (anonymous):

yes:)

jimthompson5910 (jim_thompson5910):

if the equation is \[\sqrt{x+16}-\sqrt{x}-2=0\], then the answer is x = 9 like you got katlin95, so congrats

OpenStudy (anonymous):

then yes.